OpenStudy (anonymous):

Help Please!! (: (2+y)^3 + 2y = x+27 Evaluate dy/dx at the point (2,1).

OpenStudy (anonymous):

\[(2+y)^3 + 2y = x+27\] right?

OpenStudy (anonymous):

Yes!(:

OpenStudy (anonymous):

it would probably be a lot easier to find \(\frac{dx}{dy}\) since this is actually written as a function of \(y\) but no matter take the derivative via the chain rule and get \[3(2+y)^2y'+2y'=1\]

OpenStudy (anonymous):

then solve for \(y'\) and replace \(y\) by \(1\)

OpenStudy (anonymous):

alternatively, replace \(y\) by \(1\) at the start probably easiest method

OpenStudy (anonymous):

\[3(2+y)^2y'+2y'=1\] \[3(2+1)^2y'+2y'=1\] etc

OpenStudy (anonymous):

so \[(2+1)^3+2(1)=x+27\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you are looking for \(y'\)

OpenStudy (anonymous):

wait nvm, my mistake, so how do I get y' ?

OpenStudy (anonymous):

i am sorry, i confused you with my last statement, let me start again

OpenStudy (anonymous):

you can't replace \(y\) by \(1\) at the start, you will never find \(y'\) that way what i meant is, since you have \[3(2+y)^2y'+2y'=1\] you can either a) solve for \(y'\) using algebra, then replace \(y\) by \(1\) or b) replace \(y\) by \(1\) now, then solve for \(y'\)

OpenStudy (anonymous):

So I evaluate it and get 27y' +2y'=1 29y'=1 y'= 1/29 ?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

Thank you!!!!(: @satellite73

OpenStudy (anonymous):

yw